o d.575 QUESTION9 7-13: Suppose the domestic supply (Qs) and demand (QD) for MP3 players in the United States are given by the following set of equations: Qs 25+ 10P QD 875-5P 9. in autarky (in the absence of international trade), the consumer surplus in the United States MP3 player market is e a $19,500.00 b.$33,062.50 с. $30,062.50 o d. $22,562.50 QUESTION 10 Questions 7-13: Suppose the domestic supply (Os) and demand (On) for MP3 plavers in the United States are given bv the following set of equations: Click Save and Submit to save and submit. Click Save All Answers to save all answers

when price = 60
and quantity sold = 575
Answer #1

9. The demand equation is given as: QD = 875 - 5P

inverse demand curve can be written as:

5P = 875 - QD

Or, P = 175 - 0.2QD

Therefore, the price interception of the demand curve is 175.

Equilibrium price is given as $60 and quantity is 575.

Consumer surplus = (1/2)* 575 *(175 - 60) = $ 33,062.5

SuPpl omsumer 575 2